Write an appropriate inverse variation equation if y=11 when x=3
1. y=15x
2. y=33x
3. y=33/x
4. y=11/x
Thanks so much!

Answer by stanbon(75887)

Write an appropriate inverse variation equation if y=11 when x=3
y = k/x
11 = k/3
k = 33
----
Ans:: y = 33/x
--------------------
1. y=15x
2. y=33x
3. y=33/x
4. y=11/x